OpenStudy (anonymous):

Can someone give me an example on how to factor by grouping. I know this is basic stuff, i learned it years ago, but i don't remember how to do it? i'm helping my cousin with his homework and i thought it was just factoring?

OpenStudy (anonymous):

$$8x^2-6x-2=8x^2-8x+2x-2=8x(x-1)+2(x-1)=(8x+2)(x-1)$$

OpenStudy (anonymous):

there gcf is 2, so you brake it down by 2's 8x³= 2*2*2* x*x*x 4x²= 2*2 * x*x 16x= 2*2*2*2*x each has 2*2 and *x so 4x(2x²-x-*4) The rest is in the ( )

OpenStudy (anonymous):

sorry the one i did was 8x³-4x²-16x. i got my alt codes mixed up when i typed it out the first time

OpenStudy (anonymous):

okayy. thanks!(:

OpenStudy (anonymous):

wish i could give you both medals D:

OpenStudy (anonymous):

I gave him one so it evens out